A family makes a 43.25 km trip in 5.5 hours. On the first part of the trip they crossed a lake in a canoe paddling at 12 km/h. For the rest of the trip, they hiked on a scenic trail. If their average walking speed was 5 km/h, how far did they walk?
Answer by checkley71(8403)

12T+5(5.5-T)=43.25
12T+27.5-5T=43.25
7T=43.25-27.5
7T=15.75
T=15.75/7
T=2.25 HOURS FOR THE CANOE TRIP.
5.5-2.25=3.25 HOURS FOR THE HIKE.
5*3.25=16.25 MILES WAS THE LENGTH OF THE HIKE.
PROOF
12*2.25+5*3.25=43.25
27+16.25=43.25
43.25=43.25